The formula (A∨B) ∧ (¬B ∨C) simplifies down to which of the following when using Boolean algebra?


A. A ∧ C


B. A ∧ B ∨ C


C. A ∨ C


D. B


E. A ∧ ¬B ∨ C


1
Expert's answer
2022-03-20T14:40:56-0400

Prove that none of the variants is correct:

"A:\\A=1,B=0,C=0:\left( A\lor B \right) \land \left( \lnot B\lor C \right) =1,A\land C=0\\B:\\A=1,B=1,C=0:\left( A\lor B \right) \land \left( \lnot B\lor C \right) =0,A\land B\lor C=1\\C:\\A=0,B=0,C=1:\left( A\lor B \right) \land \left( \lnot B\lor C \right) =0,A\lor C=1\\D:\\A=1,B=0,C=0:\left( A\lor B \right) \land \left( \lnot B\lor C \right) =1,B=0\\E:\\A=0,B=0,C=1:\left( A\lor B \right) \land \left( \lnot B\lor C \right) =0,A\land \lnot B\lor C=1"